A sample of gas in a balloon has an initial temperature of 33 ∘C and a volume of 1.56×103 L . If the temperature changes to 88 ∘C , and there is no change of pressure or amount of gas, what is the new volume, V2, of the gas? Express your answer with the appropriate units.
Answer #1

Ans: By Solving the problem. Initial Temperature = 33. C = 273 +33 = 306k Final Temperature - 88°C = 273 +88 = 361K And also